OpenStudy (anonymous):

Let f(x) = 2x2 - x + 5 and g(x) = 3x2 - 3x - 2. Calculate h(1), if h(x) = f(x) + g(x).

OpenStudy (angela210793):

h(x)=2x^2-x+5+3x^2-3x-2=5x^2-4x-3 h(1)=5*1-4*1-3=-2

OpenStudy (anonymous):

wait so the answer is -2?

OpenStudy (angela210793):

yea h(x)=f(x)+g(x)=2x^2-x+5+3x^2-3x-2=5x^2-4x-3 and to find h(1) all u do is replace 1 instead of x in 5x^2-4x-3

OpenStudy (anonymous):

omi god thanx u r so smart

OpenStudy (angela210793):

uw